# arXiv:2603.01692v1[cs.LG] 02 Mar 2026

### v1

@@ Abstract
LLM-based agents for machine learning engineering (MLE) predominantly rely on tree search, a form of gradient-free optimization that uses scalar validation scores to rank candidates. As LLM reasoning capabilities improve, exhaustive enumeration becomes increasingly inefficient compared to directed updates, analogous to how accurate gradients enable efficient descent over random search. We introduce Gome, an MLE agent that operationalizes gradient-based optimization. Gome maps structured diagnostic reasoning to gradient computation, success memory to momentum, and multi-trace execution to distributed optimization. Under a closed-world protocol that isolates architectural effects from external knowledge, Gome achieves a state-of-the-art 35.1% any-medal rate on MLE-Bench with a restricted 12-hour budget on a single V100 GPU. Scaling experiments across 10 models reveal a critical crossover: with weaker models, tree search retains advantages by compensating for unreliable reasoning through exhaustive exploration; as reasoning capability strengthens, gradient-based optimization progressively outperforms, with the gap widening at frontier-tier models. Given the rapid advancement of reasoning-oriented LLMs, this positions gradient-based optimization as an increasingly favorable paradigm. We release our codebase and GPT-5 traces.
Reasoning as Gradient: Scaling MLE Agents Beyond Tree Search
Yifei Zhangthanks: Equal contribution.1,2, Xu Yang11footnotemark: 11, Xiao Yang11footnotemark: 11, Bowen Xian1, Qizheng Li1,3, Shikai Fang4, Jingyuan Li1,5, Jian Wang1, Mingrui Xu4, Weiqing Liuthanks: Corresponding Author. 1, Jiang Bian1 1Microsoft Research Asia, 2Nanjing University, 3Peking University, 4Zhejiang University, 5Wuhan University, Correspondence: Weiqing.Liu@microsoft.com
@@ 1 Introduction
Automating Machine Learning Engineering (MLE) remains a long-standing challenge. Given a dataset and an evaluation metric, the objective is to autonomously complete the end-to-end development pipeline, including data preprocessing, feature engineering, model selection, and hyperparameter tuning, to deliver a high-performing predictive system. With rapid advances in code generation and complex reasoning capabilities of large language models (Jaech et al., 2024; Guo et al., 2025a), LLM-based MLE agents have recently emerged (Jiang et al., 2025; Liu et al., 2025; Nam et al., 2025), demonstrating the ability to interpret task descriptions, diagnose data characteristics, implement complete solutions, and iteratively refine them using execution feedback. To enable systematic comparison of these agents, MLE-Bench (Chan et al., 2024) provides a standardized benchmark of 75 Kaggle competitions.
@@
Current MLE agents predominantly adopt tree search, a form of gradient-free optimization. AIDE (Jiang et al., 2025) first introduced this design motivated by two practical considerations: constrained context windows that made chain-based prompting prone to prompt bloat, and multi-branch exploration to mitigate uncertainty in single-shot generations. Subsequent work largely follows this paradigm: ML-Master (Liu et al., 2025) incorporates Monte Carlo Tree Search, and AIRA (Toledo et al., 2025) proposes graph-based exploration. However, as gradient-free methods, these approaches exhibit two fundamental limitations. First, they are driven by score-based expansion: rich execution feedback (error traces, training dynamics, detailed logs) is scalarized into a single validation score that decides which node to expand, discarding the diagnostic information needed to determine how to update. This information loss becomes increasingly costly as LLM reasoning capabilities improve, since stronger models could otherwise extract precise improvement directions from such feedback. Second, these methods operate over a predefined action space: the agent selects among fixed templates, which cannot capture the effectively continuous nature of code modifications and may mismatch the specific failure modes revealed by execution.
@@
We argue that MLE inherently favors gradient-based optimization. Unlike domains punctuated by hard dead ends, MLE pipelines are typically repairable: improvements come from iteratively refining the same pipeline rather than discarding intermediate states. Moreover, the space of code modifications is effectively continuous, ranging from minor hyperparameter tweaks to architectural changes, favoring methods that propose state-conditioned updates based on execution feedback, rather than searching over a fixed action vocabulary. The key question is when such updates become preferable to enumeration. This trade-off is well understood in classical optimization: gradient-based methods dominate when gradient estimates are accurate, but degrade when noisy (Maheswaranathan et al., 2019; Asadi et al., 2024). In the MLE agent setting, the LLM’s reasoning capability determines gradient signal quality. When reasoning is weak, multi-branch enumeration hedges against unreliable proposals; as reasoning-oriented models advance, [^1] From o1-preview (Jaech et al., 2024) to DeepSeek-R1 (Guo et al., 2025a), o3 (OpenAI, 2025b), and GPT-5 (OpenAI, 2025a). the balance should shift toward gradient-based optimization: gradient-based methods should increasingly outperform tree search as reasoning capability improves (Figure 1).
@@
To test this claim, we propose Gome (Gradient-based Optimization for Machine Learning Engineering), an agent that adopts gradient-based optimization in place of enumeration-driven search. Gome uses a chain-based iterative framework where each step updates the current solution along LLM-generated improvement directions, organized into modular components that mirror classical optimizer modules (Table 2). We evaluate under a closed-world protocol to isolate architectural effects from knowledge augmentation: agents use only task-provided materials and execution feedback, without external knowledge retrieval. Under this protocol, Gome achieves a 35.1% any-medal rate on MLE-Bench with a 12-hour budget on a single V100 GPU, surpassing the previous state-of-the-art held by search-based methods. Component ablations confirm that each optimizer-inspired module contributes meaningfully. Scaling analysis, which compares Gome against tree-search baselines (including an MCTS variant sharing Gome’s codebase and previous state-of-the-art search-based methods) across 10 models ranging from non-reasoning (GPT-4o) to frontier reasoning models (GPT-5), further validates our claim: while tree-search methods show advantages with weaker reasoners, Gome’s advantage grows with model capability. Given the rapid advancement of reasoning-oriented LLMs (Sun et al., 2025; Xu et al., 2025), this positions gradient-based optimization as an increasingly favorable paradigm.
Figure 1: Comparison of search-based (gradient-free) and gradient-based optimization paradigms. (a) Tree search uses scalar scores to decide which branch to expand. (b) Gradient-based optimization uses reasoning to decide how to update; stronger reasoning yields more accurate gradient signals.
@@
In summary, this work makes the following key contributions: (1) We propose Gome, an MLE agent that adopts gradient-based optimization rather than gradient-free enumeration, establishing a functional mapping between agent components and classical optimizer modules. (2) Gome achieves state-of-the-art performance on MLE-Bench under closed-world evaluation. Scaling analysis across 10 models validates our claim: Gome’s advantage over tree search grows with model reasoning capability, confirming that stronger reasoning enables more effective gradient-based optimization. (3) We release our codebase and GPT-5 execution traces to support reproducibility and contribute to the community.
@@ 2 Related Work
Current MLE agents explore various search topologies: tree search (AIDE (Jiang et al., 2025), ML-Master (Liu et al., 2025), KompeteAI (Kulibaba et al., 2025)), graph-based frameworks (AIRA (Toledo et al., 2025), InternAgent-MLE (InternAgent-Team, 2025)), and evolutionary methods (FM Agent (Li et al., 2025)). Some approaches augment search with external knowledge retrieval from Kaggle notebooks, arXiv papers, or expert knowledge bases. MLE-STAR (Nam et al., 2025) adopts a chain structure with web retrieval and targeted code-block refinement; however, its iteration remains score-driven, using ablation-based scoring to select which block to refine rather than reasoning about how to update (see Appendix D.2). We summarize key design choices in Table 1. Despite structural differences, these methods use execution feedback primarily for ranking: scalar validation scores select which branch to expand or which candidate to retain. In contrast, Gome integrates scalar scores with execution feedback for update: structured reasoning over these signals generates improvement hypotheses that directly guide solution modification.
@@
Table 1: Comparison of LLM-based MLE agents.
@@
A growing line of work operationalizes LLM reasoning as an optimization signal, using task outcomes to iteratively refine artifacts rather than exhaustively enumerating candidates (Yang et al., 2023; Shinn et al., 2023; Zhang et al., 2025). In prompt optimization, LLMs propose edits conditioned on performance feedback, with methods ranging from discrete optimization procedures (Pryzant et al., 2023; Cui et al., 2024) to population-based evolution (Guo et al., 2025b). In agentic settings, trial outcomes are converted into self-reflections that guide behavior updates (Madaan et al., 2023), while textual “gradients” formalize refinement as structured edits (Yuksekgonul et al., 2024; Cheng et al., 2024). While this paradigm has been widely adopted in prompt tuning and general-purpose agent tasks, it remains underexplored in MLE, where existing methods still rely on score-based search (Table 1). Gome bridges this gap, instantiating reasoning-as-optimization for MLE with structured feedback as gradient signal for code updates.
@@ 3 Gome
Figure 2: Overview of the Gome framework. Multiple traces optimize in parallel, synchronizing through a global shared success memory $\mathcal{M}$. Each trace iteratively executes solutions, validates improvements, updates shared memory, and reasons over local and global feedback to generate the next hypothesis. Right panels illustrate the gradient-based optimization analogy and key module details.
Our goal is to find an optimal solution within the code space under resource constraints:
@@
Existing MLE agents typically organize $\mathcal{S}$ as a tree and use scalar scores to decide which node to expand, a form of gradient-free optimization. We propose Gome, which instead focuses on how to update. Inspired by gradient-based optimization, Gome extracts directional improvement signals from structured feedback via LLM reasoning. While true gradients are undefined in discrete code space, LLM reasoning can serve a functionally analogous role: analyzing execution results to determine not just whether a solution improved, but why it improved and what to change next. Table 2 summarizes this analogy.
Figure 2 illustrates the framework. Given task specification $\mathcal{T}$ and dataset $\mathcal{D}$, Gome initializes $N$ parallel optimization traces, each beginning with a distinct hypothesis (i.e., an improvement direction for the ML pipeline) through forced diversification (§3.5). Each trace maintains its own local best solution $s^{*(i)}$ and experiment history $\mathcal{H}^{(i)}$, while contributing to and drawing from a shared success memory $\mathcal{M}$. Each iteration proceeds through four stages: (1) execute the current solution to obtain execution feedback (§3.1); (2) apply hierarchical validation to produce diagnostic reasoning and accept decision, forming structured feedback $f^{(i)}_{t}$3.2); (3) if accepted, contribute the successful hypothesis to $\mathcal{M}$3.3); (4) generate the next hypothesis $\eta^{(i)}_{t+1}$ by combining local feedback with shared memory (§3.4). This interplay between local execution and shared learning enables collaborative optimization across traces (§3.5). The complete algorithm is provided in Appendix C.
@@
Momentum Success Memory Acceleration via global proven patterns
@@
Table 2: Analogy between Gome components and gradient-based optimization.
@@ 3.2 Hierarchical Validation
A critical challenge in LLM-based optimization is distinguishing genuine improvements from deceptive ones. Solutions may achieve high validation scores through shortcuts such as data leakage or metric gaming. Based on execution feedback, hierarchical validation analyzes the solution and determines acceptance:
$(d^{(i)}_{t},\texttt{reason}_{t})=\texttt{Validate}(s^{(i)}_{t},\texttt{perf}_{t},\texttt{trace}_{t}),$
where $d^{(i)}_{t}\in\{0,1\}$ is the accept decision and $\texttt{reason}_{t}$ provides diagnostic analysis. Together with execution feedback, this forms the complete structured feedback:
$f^{(i)}_{t}=(\texttt{perf}_{t},\texttt{trace}_{t},\texttt{reason}_{t}).$ (4)
Three-Stage Validation.
The validation implements three sequential stages: (1) Format correctness verifies output files exist and match the required schema via rule-based checks. (2) Evaluation alignment detects data leakage and overfitting risks via LLM analysis, identifying violations such as accessing test labels during training or using future information in time-series splits. (3) Comprehensive analysis verifies whether the hypothesis achieved its intended effect and analyzes code quality when $h(s^{(i)}_{t})\gtrsim h(s^{*(i)})$; the final accept decision is made by LLM based on this reasoning. The diagnostic output $\texttt{reason}_{t}$ captures the analysis from whichever stage determined the outcome, informing subsequent hypothesis generation.
@@ 3.3 Success Memory
When a solution passes validation ($d^{(i)}_{t}=1$), the improvement hypothesis $\eta^{(i)}_{t}$ that generated this solution, along with its feedback, are added to the shared success memory:
$\mathcal{M}_{t+1}=\begin{cases}\mathcal{M}_{t}\cup\{(\eta^{(i)}_{t},f^{(i)}_{t},\Delta h^{(i)}_{t})\}&\text{if }d^{(i)}_{t}=1\\ \mathcal{M}_{t}&\text{otherwise}\end{cases}$
where $\eta^{(i)}_{t}$ is the hypothesis that led to solution $s^{(i)}_{t}$, $f^{(i)}_{t}$ is the structured feedback (§3.2), and $\Delta h^{(i)}_{t}=h^{(i)}_{t}-h^{*(i)}$ records the score difference (which may be negative if accepted for other merits such as code quality). Each memory entry thus contains the complete context: what was tried (hypothesis), what happened (feedback), and how much it helped (score delta). $\mathcal{M}$ differs from local history $\mathcal{H}^{(i)}$ in two ways: it stores only verified successes (not all experiments), and aggregates across all traces (not just one). In single-trace settings, $\mathcal{M}$ serves as a curated subset of $\mathcal{H}^{(i)}$. Functionally analogous to momentum (Table 2), success memory biases future updates toward proven directions; how this memory is leveraged across traces is detailed in §3.5.

### v2

# Reasoning as Gradient: Scaling MLE Agents Beyond Tree Search

@@ Abstract
Abstract: LLM-based agents for machine learning engineering (MLE) predominantly rely on tree search, a form of gradient-free optimization that uses scalar validation scores to rank candidates. As LLM reasoning capabilities improve, exhaustive enumeration becomes increasingly inefficient compared to directed updates, analogous to how accurate gradients enable efficient descent over random search. We introduce Gome, an MLE agent that operationalizes gradient-based optimization. Gome maps structured diagnostic reasoning to gradient computation, success memory to momentum, and multi-trace execution to distributed optimization. Under a closed-world protocol that isolates architectural effects from external knowledge, Gome achieves a state-of-the-art 35.1% any-medal rate on MLE-Bench with a restricted 12-hour budget on a single V100 GPU. Scaling experiments across 10 models reveal a critical crossover: with weaker models, tree search retains advantages by compensating for unreliable reasoning through exhaustive exploration; as reasoning capability strengthens, gradient-based optimization progressively outperforms, with the gap widening at frontier-tier models. Given the rapid advancement of reasoning-oriented LLMs, this positions gradient-based optimization as an increasingly favorable paradigm. We release our codebase and GPT-5 traces at https://github.com/microsoft/RD-Agent.
11footnotetext: Equal contribution.22footnotetext: Corresponding Author.
@@ 1 Introduction
Automating Machine Learning Engineering (MLE) remains a long-standing challenge. Given a dataset and an evaluation metric, the objective is to autonomously complete the end-to-end development pipeline, including data preprocessing, feature engineering, model selection, and hyperparameter tuning, to deliver a high-performing predictive system. With rapid advances in code generation and complex reasoning capabilities of large language models [jaech2024openai, guo2025deepseek], LLM-based MLE agents have recently emerged [aide2025, liu2025ml, nam2025mle], demonstrating the ability to interpret task descriptions, diagnose data characteristics, implement complete solutions, and iteratively refine them using execution feedback. To enable systematic comparison of these agents, MLE-Bench [chan2024mle-bench] provides a standardized benchmark of 75 Kaggle competitions.
@@
Current MLE agents predominantly adopt search-based exploration. AIDE [aide2025] first introduced this design motivated by two practical considerations: constrained context windows that made chain-based prompting prone to prompt bloat, and multi-branch exploration to mitigate uncertainty in single-shot generations. Subsequent work largely follows this paradigm: ML-Master [liu2025ml] incorporates Monte Carlo Tree Search, and AIRA [toledo2025ai] proposes graph-based exploration. Importantly, these methods use directed topologies (e.g., UCT/PUCT-style node selection with value backpropagation). Our characterization of “gradient-free” is therefore about the optimization signal, not the search topology: improvement is still primarily selected by scalar rewards/scores rather than by explicitly reasoning about update direction from rich execution feedback. Under this signal-level view, prior search-based agents exhibit two limitations. First, they are driven by score-centric selection: rich execution feedback (error traces, training dynamics, detailed logs) is compressed into scalar rewards that mainly decide which node/branch survives, discarding diagnostic information needed to determine how to update. This information loss becomes increasingly costly as LLM reasoning capabilities improve, since stronger models could otherwise extract precise improvement directions from such feedback. Second, these methods operate over a predefined action space: the agent selects among fixed templates, which cannot capture the effectively continuous nature of code modifications and may mismatch the specific failure modes revealed by execution.
@@
We argue that MLE inherently favors gradient-based optimization. Unlike domains punctuated by hard dead ends, MLE pipelines are typically repairable: improvements come from iteratively refining the same pipeline rather than discarding intermediate states. Moreover, the space of code modifications is effectively continuous, ranging from minor hyperparameter tweaks to architectural changes, favoring methods that propose state-conditioned updates based on execution feedback, rather than searching over a fixed action vocabulary. The key question is when such updates become preferable to enumeration. This trade-off is well understood in classical optimization: gradient-based methods dominate when gradient estimates are accurate, but degrade when noisy [maheswaranathan2019guided, asadi2024comparative]. In the MLE agent setting, the LLM’s reasoning capability determines gradient signal quality. When reasoning is weak, multi-branch enumeration hedges against unreliable proposals; as reasoning-oriented models advance, [^1] From o1-preview [jaech2024openai] to DeepSeek-R1 [guo2025deepseek], o3 [openai_introducing_o3_2025], and GPT-5 [openai_introducing_gpt-5_2025]. the balance should shift toward gradient-based optimization: gradient-based methods should increasingly outperform tree search as reasoning capability improves (Figure 1).
@@
To test this claim, we propose Gome (Gradient-based Optimization for Machine Learning Engineering), an agent that adopts gradient-based optimization in place of enumeration-driven search. Gome uses a chain-based iterative framework where each step updates the current solution along LLM-generated improvement directions, organized into modular components that mirror classical optimizer modules (Table 2). We evaluate under a closed-world protocol to isolate architectural effects from knowledge augmentation: agents use only task-provided materials and execution feedback, without external knowledge retrieval. Under this protocol, Gome achieves a 35.1% any-medal rate on MLE-Bench with a 12-hour budget on a single V100 GPU, surpassing the previous state-of-the-art held by search-based methods. Component ablations confirm that each optimizer-inspired module contributes meaningfully. Scaling analysis, which compares Gome against tree-search baselines (including an MCTS variant sharing Gome’s codebase and previous state-of-the-art search-based methods) across 10 models ranging from non-reasoning (GPT-4o) to frontier reasoning models (GPT-5), further validates our claim: while tree-search methods show advantages with weaker reasoners, Gome’s advantage grows with model capability. Given the rapid advancement of reasoning-oriented LLMs [sun2025survey, xu2025towards], this positions gradient-based optimization as an increasingly favorable paradigm.
Figure 1: Two paradigms for utilizing execution feedback. (a) Score-centric: feedback is compressed into scalar rewards that guide node/branch selection (which may itself be directed, e.g., via UCT/PUCT). (b) Gradient-based: structured reasoning over execution feedback determines how to update the current solution; stronger reasoning yields more accurate signals.
@@
In summary, this work makes the following key contributions: (1) We propose Gome, an MLE agent that adopts gradient-based optimization rather than score-centric candidate ranking, establishing a functional mapping between agent components and classical optimizer modules. (2) Gome achieves state-of-the-art performance on MLE-Bench under closed-world evaluation. Scaling analysis across 10 models validates our claim: Gome’s advantage over tree search grows with model reasoning capability, confirming that stronger reasoning enables more effective gradient-based optimization. (3) We release our codebase and GPT-5 execution traces to support reproducibility and contribute to the community.
@@ 2 Related Work
Current MLE agents explore various search topologies: tree search (AIDE [aide2025], ML-Master [liu2025ml], KompeteAI [kulibaba2025kompeteai]), graph-based frameworks (AIRA [toledo2025ai], InternAgent-MLE [internagent-mle]), and evolutionary methods (FM Agent [li2025fm]). Some approaches augment search with external knowledge retrieval from Kaggle notebooks, arXiv papers, or expert knowledge bases. MLE-STAR [nam2025mle] adopts a chain structure with web retrieval and targeted code-block refinement; however, its iteration remains score-driven, using ablation-based scoring to select which block to refine rather than reasoning about how to update (see Appendix D.3). We summarize key design choices in Table 1. Notably, AIRA formalizes research agents as $(\mathcal{F},\pi_{\text{sel}},\mathcal{O},\pi_{\text{op}},\tau)$ and shows that, under the AIDE operator set, changing the search policy alone yields only limited gains, suggesting that operator design can be a major bottleneck beyond search topology [toledo2025ai]. InternAgent-MLE [internagent-mle] further extends tree search to graph search with cross-branch trajectory reuse, yet the optimization signal remains score-driven. These examples illustrate a common pattern: topologies grow increasingly directed, but feedback utilization stays score-centric. In contrast, Gome integrates scalar scores with execution feedback for update: structured reasoning over these signals generates improvement hypotheses that directly guide solution modification.
@@
Table 1: Comparison of LLM-based MLE agents. Feedback Role indicates how execution feedback drives the primary optimization loop: Ranking denotes score-centric selection among candidates; Update denotes diagnostic reasoning over structured feedback to generate state-conditioned code modifications.
@@
A growing line of work operationalizes LLM reasoning as an optimization signal, using task outcomes to iteratively refine artifacts rather than exhaustively enumerating candidates [yang2023large, shinn2023reflexion, zhang2025process]. In prompt optimization, LLMs propose edits conditioned on performance feedback, with methods ranging from discrete optimization procedures [pryzant2023automatic, cui2024introducing] to population-based evolution [guo2025evoprompt]. In agentic settings, trial outcomes are converted into self-reflections that guide behavior updates [madaan2023self], while textual “gradients” formalize refinement as structured edits [yuksekgonul2024textgrad, cheng2024trace]. While this paradigm has been widely adopted in prompt tuning and general-purpose agent tasks, it remains underexplored in MLE, where existing methods still rely on score-based search (Table 1). Gome bridges this gap, instantiating reasoning-as-optimization for MLE with structured feedback as the gradient signal for code updates.
@@ 3 Gome
We formulate MLE tasks as finding an optimal solution within the code space under resource constraints:
@@
Existing MLE agents typically organize $\mathcal{S}$ as a tree or graph and use scalar scores to guide expansion. We propose Gome, which instead focuses on how to update. Inspired by gradient-based optimization, Gome extracts directional improvement signals from structured feedback via LLM reasoning. While true gradients are undefined in discrete code space, LLM reasoning serves a functionally analogous role: analyzing execution results to determine not just whether a solution improved, but why it improved and what to change next. Table 2 summarizes this analogy.
Figure 2 illustrates the framework. Gome initializes $N$ parallel optimization traces, each beginning with a distinct hypothesis (i.e., an improvement direction for the ML pipeline) through forced diversification (§3.5). Each trace maintains local best solution $s^{*(i)}$ and experiment history $\mathcal{H}^{(i)}$, while synchronizing via a shared success memory $\mathcal{M}$. Each iteration proceeds through four stages: (1) Execution: run current solution and collect feedback (§3.1); (2) Validation: apply hierarchical checks to form structured feedback $f_{t}^{(i)}$3.2); (3) Memory update: contribute successful hypotheses to $\mathcal{M}$ (Momentum accumulation) (§3.3); (4) Reasoning: generate the next hypothesis $\eta^{(i)}_{t+1}$ by combining local feedback with shared memory (Gradient computation) (§3.4). This local-global interplay enables collaborative optimization (§3.5), with full algorithmic details in Appendix C.
@@
Momentum Success Memory Acceleration via proven patterns
@@
Table 2: Analogy between Gome and gradient-based optimization.
Figure 2: Overview of the Gome framework. Multiple traces optimize in parallel, synchronizing through a global shared success memory $\mathcal{M}$. Each trace iteratively executes solutions, validates improvements, updates shared memory, and reasons over local-global feedback to generate the next hypothesis. Right panels illustrate the gradient-based optimization analogy and key module details.
@@ 3.2 Hierarchical Validation
A critical challenge in LLM-based optimization is distinguishing genuine improvements from deceptive ones. Solutions may achieve high validation scores through shortcuts such as data leakage or metric gaming. Based on execution feedback, hierarchical validation determines acceptance decision $d^{(i)}_{t}\in\{0,1\}$ and diagnostic $\texttt{reason}_{t}$:
$(d^{(i)}_{t},\texttt{reason}_{t})=\texttt{Validate}(s^{(i)}_{t},\texttt{perf}_{t},\texttt{trace}_{t}),$ (3)
forming the structured feedback $f^{(i)}_{t}=(\texttt{perf}_{t},\texttt{trace}_{t},\texttt{reason}_{t})$. The validation process enforces three sequential stages: (1) Format correctness: verifies output schema via rule-based checks; (2) Evaluation alignment: detects data leakage and overfitting risks via LLM analysis, e.g., accessing test labels or future information; (3) Comprehensive analysis: verifies whether the hypothesis achieved its intended effect and assesses code quality. The LLM makes the final accept decision $d^{(i)}_{t}$ based on this reasoning when $h(s^{(i)}_{t})\gtrsim h(s^{*(i)})$, with $\texttt{reason}_{t}$ capturing the analysis from whichever stage determined the outcome, informing subsequent hypothesis generation. Empirically, hierarchical validation detects 66.7% of deceptive overfitting attempts, compared to 0% for purely score-driven baselines (Appendix D.1).
@@ 3.3 Success Memory
Upon validation, the hypothesis and its feedback are committed to the shared success memory $\mathcal{M}$:
$\mathcal{M}_{t+1}=\begin{cases}\mathcal{M}_{t}\cup\{(\eta^{(i)}_{t},f^{(i)}_{t},\Delta h^{(i)}_{t})\}&\text{if }d^{(i)}_{t}=1\\ \mathcal{M}_{t}&\text{otherwise}\end{cases}$ (4)
where $\eta^{(i)}_{t}$ is the hypothesis that led to solution $s^{(i)}_{t}$, $f^{(i)}_{t}$ is the structured feedback (§3.2), and $\Delta h^{(i)}_{t}=h^{(i)}_{t}-h^{*(i)}$ records the score difference (may be negative if accepted for code quality merits). Each memory entry contains the complete context: what was tried (hypothesis), what happened (feedback), and how much it helped (score delta). $\mathcal{M}$ is distinct from the raw local history $\mathcal{H}^{(i)}$: in single-trace settings, it serves as a curated subset of verified successes; in multi-trace settings, it naturally extends to a global repository aggregating discoveries across all workers. Functionally analogous to momentum, this global memory biases future updates toward proven directions.